Calculate the Energy (in Joules) of a photon emitted that has a wavelength of 336nm.
1
Expert's answer
2019-10-07T06:42:28-0400

E = hv = h(c/λ) = 6.626*10-34(3*108/336*10-9 = 5.916*10-19J